# Agnostics: Learning to Code in Any Programming Language via Reinforcement with a Universal Learning Environment

### v1

@@ Abstract
Large language models (LLMs) already excel at writing code in high-resource languages such as Python and JavaScript, yet stumble on low-resource languages that remain essential to science and engineering. Besides the obvious shortage of pre-training data, post-training itself is a bottleneck: every new language seems to require new datasets, test harnesses, and reinforcement-learning (RL) infrastructure.
@@
Applied to five low-resource languages—Lua, Julia, R, OCaml, and Fortran—Agnostics (1) improves Qwen-3 4B to performance that rivals other 16B–70B open-weight models; (2) scales cleanly to larger and diverse model families (Qwen-3 8B, DeepSeek Coder 6.7B Instruct, Phi 4 Mini); and (3) for ${\leq}16$B parameter models, sets new state-of-the-art pass@1 results on MultiPL-E and a new multi-language version LiveCodeBench that we introduce.
We will release the language-agnostic training datasets (Ag-MBPP-X, Ag-Codeforces-X, Ag-LiveCodeBench-X), training code, and ready-to-use configurations, making RL post-training in any programming language as simple as editing a short YAML file.
@@ 1 Introduction
Figure 1: Overview: Agnostics Data Preparation and Training. (1) We reformulate existing coding datasets to our format. (2) We adapt the language-agnostic datasets to a particular programming language. (3, 4) We reinforce coding via Group Relative Policy Optimization (Shao et al. 2024; DeepSeek-AI et al. 2025), verifying the programs in our code execution sandbox.
@@
Large language models (LLMs) are remarkably good at programming tasks, especially when codinng in high-resource programming languages such as Python and JavaScript. Their proficiency in low-resource programming languages, such as Fortran, Julia, and others, is significantly more limited. This gap appears both on benchmarks (Cassano et al. 2023) and in popular discourse. However, many low-resource languages are widely used in computational science (e.g., Fortran), medicine (e.g., Mumps), engineering (e.g., Matlab), and others sectors. For programmers in these sectors to be able to take full advantage of LLMs, we need methods to make LLMs better at the languages that they use.
@@
The capability gap between high-resource and low-resource languages occurs for two reasons. The first one is that there is vastly more training data for some programming languages. For example, The Stack V2 (Lozhkov et al. 2024a), the largest public training corpus of code, has ${\approx}200$GB of Python but only ${\approx}2$GB of Julia and Fortran. Thus pretraining on code makes models significantly better at Python. A more subtle reason is the availability of post-training datasets and techniques. Contemporary LLMs are developed with an extensive post-training process that relies on (a) high-quality curated data for supervised fine-tuning, and (b) carefully designed environments for reinforcement learning, which must be able to execute and verify model-generated solutions. Both of these require significant human expertise, which is hard to find for low-resource languages.
@@
This paper aims to dramatically facilitate post-training LLMs on low-resource programming languages, working towards closing the resource gap. Our key idea is that for a significant class of programming tasks, correctness can be specified as a property not of individual functions or code snippets, but of the entire program’s observable behavior (e.g., I/O). Furthermore, if its correctness can be tested with another program, we can use such a verifier together with appropriate problems and test cases to make a universal reinforcement learning environment which can be adapted to any programming language. In fact, the language used to implement the verifier and the one being learned are entirely independent. This approach matches how some existing post-training datasets are formulated (even though they are intended for Python/C++). We can also use LLMs to reformulate language-specific datasets into this format.
@@
Based on this insight, our approach, Agnostics, works as follows (Figure 1). 1) We use an LLM to reformulate language-specific datasets into our standard language-agnostic format. 2) We generate prompts and a configured verifier targeting a particular language, with a small (4-5 line) configuration. 3) We apply reinforcement learning with verified rewards (RLVR) using a robust, language-agnostic execution sandbox that we develop. 4) The result is a model specialized to the target language. Agnostics particularly excels at finetuning models for low-resource languages, as it does not rely on high-quality datasets specific to a particular language.
@@
    Agnostics, a post-training pipeline for coding in arbitrary programming languages, including low-resource ones;
@@
    The best-performing open-weights models for Lua, R, Julia, OCaml and Fortran at the ${\leq}16$B parameter scale;
@@
    3 Agnostics datasets Ag-MBPP-X, Ag-Codeforces-X, and Ag-LiveCodeBench-X, based on MBPP (Austin et al. 2021), Open-R1 Codeforces (Penedo et al. 2025) and LiveCodeBench (Jain et al. 2024a) respectively.
@@
    A small and highly-optimized Agnostics training framework, implementing GRPO and including a parallel code execution sandbox, sampling, rewards computation, and model back-propagation.
Artifacts and code necessary to verify the results we present are listed at agnostics.abgru.me. The training framework will be publicly released together with the formal publication of this work.
@@ 2 Background and Related Work
Figure 2: High-resource languages (left) dominate the Stack v2. We are interested in low-resource languages (right).
Data Scarcity for Low-Resource Languages
LLM have been pretraining on code for several years, both because LLMs are widely used for practical programming tasks, and because pretraining on code improves their general reasoning abilities (Ma et al. 2023). In addition to general-purpose LLMs, there are also code-specialized models that are either trained exclusively on code (e.g., Xu et al. (2022); Allal et al. (2023)) or trained on code starting from a general-purpose model checkpoint (e.g., Rozière et al. (2024)).
However, the publicly available pretraining data for code is heavily skewed toward a handful of programming languages. For example, consider The Stack V2 (Lozhkov et al. 2024b), which is the largest public code pretraining dataset, with code from GitHub and dozens of other sources. The Stack V2 dominated by a handful of programming languages: just 10 of 619 languages account for over 90% of the dataset. We are interested in developing models for low-resource languages that each account for less than 0.5% of the publicly available code (figure 2).
We can imagine working around this data scarcity in a few ways. However, previous work shows that up-sampling low-resource languages during pretraining only leads to small benchmark improvements (Orlanski et al. 2023), while fine-tuning on the pre-training data for low-resource languages has negligible impact (Cassano et al. 2024). Thus, it is not clear how to make further gains from existing natural data.
@@
If natural data is not available for a task, it is possible to use LLMs to generate synthetic fine-tuning data (Wang et al. 2023), and there are many techniques for building code-centric supervised fine-tuning datasets (e.g. Luo et al. (2023); Wei et al. (2024c, b)) that work remarkably for Python. Although these approaches could in principle be applied to any programming language, Cassano et al. (2024) show that without distillation or verification (e.g., see Hu et al. (2025); Wei et al. (2024a)), synthetic tasks and code for low-resource programming languages are low-quality, and models fine-tuned on them perform poorly.
MultiPL-T (Cassano et al. 2024), similar to TransCoder-ST (Roziere et al. 2021) and CMTrans (Xie et al. 2024), couples synthetic data generation with verification using rejection sampling: it generates up $n$ candidate programs in a target, low-resource language and only fine-tune models on generations that pass hidden unit tests that it translates from Python. However, the MultiPL-T approach has two significant limitations. (1) For the verifier to not reject all samples, the model must be able to generate a working program within $n$ attempts. In MultiPL-T, ${\approx}30\%$ of prompts produce a working program for $n\in\{50,100\}$ attempts, and the rest are discarded. We train on much harder problems, and estimate that rejection sampling would require an order of magnitude more resources to achieve a comparable acceptance rate (§ A). (2) For each low-resource language of interest, MultiPL-T requires the user to write a little compiler to translate test cases and function signatures from Python to the target language. MultiPL-T only supports a limited set of built-in Python types (e.g., no classes) and it dictates that all Python types and values must faithfully map to the target language. However, depending on the problem and language, the natural way to represent data may not map cleanly to Python types. This can lead to peculiar, unidiomatic translations that require deep language expertise to get right. The Agnostics approach is significantly easier to use than MultiPL-T, and only requires the user to know how to compile and run a program in the target language from the shell.
@@
DeepSeek R1 (DeepSeek-AI et al. 2025) popularized RL on LLMs with rule-based rewards, instead of learned reward models. R1 reports applying RL to coding tasks but does not disclose further dataset details. A number of papers apply RL to the NL to code task (Zeng et al. 2025; Gehring et al. 2024; Jain et al. 2025). These techniques, which target Python, show that reinforcement learning can improve LLM capabilities beyond what is possible with just supervised fine-tuning.
However, the key benefit of reinforcement learning is that it can train a model to do tasks for which high-quality data for supervised fine-tuning is unavailable. There are recent examples of using RL for code optimization (Du et al. 2025; Nichols et al. 2024), resolve GitHub, issue resolution (Wei et al. 2025), and iterative development (Zhou et al. 2025). These papers target tasks in high-resource languages—C++, Java, and Python—whereas Agnostics targets several low-resource languages.
@@ 3 The Agnostics Approach
Our approach comprises (1) a data preparation stage which reformulates language-specific programming tasks to be language-agnostic, and retargets language-agnostic datasets to a programming language of interest (1, 2 in figure 1); and (2) the training stage which uses the GRPO algorithm with an efficient, language-agnostic verification framework (3, 4 in figure 1). Our central idea is to have tasks ask for programs which behave in a particular way. The test cases are samples of this behavior, and a verifier program can check if a solution behaves according to the sample. In this paper, we limited ourselves to working with tasks asking for programs which read data from the standard input, compute a unique answer, and write it to the standard output. Hence, the datasets we prepared share the same verifier.
# Write a python function to identify non-
# prime numbers.
@@
(a) A task prompt and associated tests (in gray) from MBPP.
Instruction: Given an integer NN ($N\geq 2$), determine whether it is a non-prime number. Output ‘True’ if the number is non-prime, ‘False’ otherwise. Input format: a single integer NN ($N\geq 2$). Output format: a single line containing ‘True’ or ‘False’.
@@ 3.1 Dataset Preparation
Some datasets, like Open-R1 Codeforces (Penedo et al. 2025), already define tasks in the desired I/O style. More commonly, however, code datasets provide a set of unit tests. Figure 3(a) shows a representative item from MBPP: it has a natural language problem description and a Python function signature that comprise the prompt, and a suite of tests used to test model-generated code. These datasets can be easily translated into the I/O format.
To make such problems language‑agnostic and compatible with our verifier, we prompt an LLM to reformulate each task so that the program communicates exclusively via plain‑text standard in and standard out. We ask the model to spell out concrete I/O conventions—number of decimal places, newline versus comma separators, ordering of values, and so on—so that the expected behavior is unambiguous. Figure 3(b) shows the reformulated example. § B has the instruction we use to reformulate MBPP; other datasets might require small changes to the prompt.
@@ 3.2 Programming Language Preparation
Use R version 4.
Use readLines(con = file("stdin"))‘ to
read input from stdin. Optionally, use
the n argument to read the first n
lines. For example:
@@
print(n) # print the first line of input
Also, use cat to print output to stdout.
For example:
‘‘‘r
cat(n)
Please do not use print to print output.
Figure 4: An Agnostics configuration snippet for R.
To prepare a new language, we author a small configuration file that serves two purposes. First, it defines a prompt prefix (prepended to each problem by the trainer) which instructs the model to produce code in the target language. Second, the configuration file specifies the shell commands to install the language toolchain and run code. Figure 4 shows the configuration for the R language, with an atypically large prompt due to quirks of its I/O APIs.111 There are three ways to run R, three I/O APIs, and only one portable way to read from standard in.
A good prefix should help avoid the most common errors, but determining which errors are common can be time-consuming. In our experience, if the model starts from ${\geq}5\%$ accuracy, then the prefix only needs to say that the solution must be in language $L$. (This varies with GRPO group size, see § 3.3.) On the other hand, when starting from near-zero accuracy, a prefix like in figure 4 can help prevent common mistakes. For OCaml and Fortran, we asked a capable LLM (OpenAI o3) for advice based on several faulty generations.
We used the resulting instructions verbatim (§ C), but only for training and not to evaluate models on OCaml or Fortran.
@@ 3.3 Trainer and Code Execution
The Agnostics trainer uses the Group-Relative Policy Optimization (GRPO) reinforcement learning algorithm (Shao et al. 2024), with verifiable rewards (DeepSeek-AI et al. 2025), and some common tweaks to further improve its efficiency (Yu et al. 2025). We couple the algorithm with a language-agnostic code execution framework that is designed to be robust and efficient.
@@
The trainer instantiates the GRPO algorithm as follows. Let $(x,\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K})\sim\mathcal{D}$ be a dataset of language-agnostics tasks, where $x$ is the task prompt and $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$ is the set of I/O examples. Let $L$ be a language configuration $L$ (e.g., figure 4). From the behavior policy $\pi_{\theta_{\text{old}}}$, we sample a group $G$ of candidate responses
$\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(y_{i}\mid L.\mathsf{prompt},x).$
We assign each candidate a reward $R_{i}$, with $R_{i}=1$ if the execution environment (described later) verifies that the extracted program’s behavior is consistent with the I/O examples $(\mathit{in}_{k},\mathit{out}_{k})$, and $R_{i}=0$ otherwise.
We turn group rewards into sequence-level advantages
$\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}(\{R_{j}\}_{j=1}^{G})}{\operatorname{std}(\{R_{j}\}_{j=1}^{G})}.$
We then update the policy with the objective
$\mathcal{L}_{\text{GRPO}}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Bigl{(}r_{i,t}(\theta)\hat{A}_{i},c_{i,t}(\theta)\hat{A}_{i}\Bigr{)},$
where
$c_{i,t}(\theta)$ $=\text{clip}\bigl{(}r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\bigr{)},$
$r_{i,t}(\theta)$ $=\frac{\pi_{\theta}(y_{i,t}\mid L.\mathsf{prompt},x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid L.\mathsf{prompt},x,y_{i,<t})}.$
We omit the KL-divergence term, similar to Yu et al. (2025).
We considered and decided against a reward for a partially-correct answer. We tried to reward the model for code which runs without errors but produces wrong output or for code which only passes the public tests. In both cases the models were very likely to learn how to exploit the reward, e.g., by producing empty programs or by hard-coding the public tests (and claiming to produce a “draft answer”).
@@
Our verifier is a language-agnostic code execution sandbox which (1) extracts a program from each candidate; (2) compiles it if needed; and (3) tests it on the I/O examples $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$.
To extract the code, we instruct the model to put it in a Markdown block, which all major instruction-tuned models do by default. Since we rely on the native format of the model, we do not need to train the model with a format reward. This guarantees that the increases in rewards that we see during training are genuine improvements and not merely the result of the model learning to format correctly.
For each language, we build and cache an OCI container using the configuration $L$. To build the container, we install the language compiler and runtime (the script $L.\texttt{install}$), and include a generic execution harness which runs and tests candidate programs. The execution harness runs continuously in the container, waiting to receive a triple with the candidate program, the set of input/output examples, and timeouts. The harness (1) writes the program to disk (to $L.\texttt{filename}$), (2) compiles it if needed ($L.\texttt{compile}$), (3) runs it on each received input ($L.\texttt{execute}$), and verified that it produces the expected output. The harness imposes timeouts on the compilation step and each execution, and returns reward 0 on any timeout or failed verification.
It is important to have timeouts for both compilation and execution. This prevents pathologies such as unbounded macro expansion in Julia (caught by the compile timeout) and infinite loops (caught by the execution timeout). Using containers also allows us to limit CPU, memory, and filesystem usage; no elevated privileges are granted to the generated program. Although the current datasets only specify tasks by standard I/O, the same sandbox can safely accommodate problems that read/write from network or disk.
@@
Finally, to improve compile times, our execution environment mounts a RAM disk in each container. Compilation may be slow due to creating many intermediate files, and indeed some large C++ projects, e.g., Firefox, recommend using a RAM disk to speed up their builds (Firefox 2025).
@@
We implement the trainer and execution environment with Ray (Moritz et al. 2018), which facilitates multiprocessing and distributed computing. In particular, the framework allows distributing the training over a network of heterogeneous nodes, which allows running the trainer on a node specialized for GPU work and the execution environment on a node specialized for CPU work. Ray also lets us easily have two separate inter-communicating processes, one for group generation and another for verification and loss computation, simply by defining two actor classes. Running the two concurrently significantly speeds up training, as we found that they take a roughly comparable amount of time. The execution environment is also implemented as an actor, with one Python asyncio task per container. Using tasks instead of actors helps lower how much inter-process communication and data copying is needed.
@@
For our training, we use the AdamW optimizer (Loshchilov and Hutter 2019) with a learning rate of $5\times 10^{-6}$ and a cosine decay schedule with a warmup of 0.1 epochs. We process 4 prompts in each batch, with group size 32 per prompt. We generate with temperature 0.7. We train hybrid reasoning models with reasoning disabled. (We still often observe generations with reasoning-like text, either before the answer or in comments.)

### v2

# Agnostics: Learning to Synthesize Code in Any Programming Language with a Universal Reinforcement Learning Environment

@@ Abstract
Large language models (LLMs) excel at writing code in high-resource languages such as Python and JavaScript, yet stumble on low-resource languages that remain essential to science and engineering. Besides the obvious shortage of pre-training data, post-training itself is a bottleneck: every new language seems to require new datasets, test harnesses, and reinforcement-learning (RL) infrastructure.
@@
Applied to five low-resource languages—Lua, Julia, R, OCaml, and Fortran—Agnostics (1) improves Qwen-3 4B to performance rivaling other 16B–70B open-weight models; (2) scales to larger and diverse model families (Qwen-3 8B, DeepSeek Coder 6.7B Instruct, SmolLM 3, Phi 4 Mini); and (3) for ${\leq}16$B parameter models, sets new state-of-the-art pass@1 results on MultiPL-E and a new multi-language version of LiveCodeBench which we introduce.
We release the language-agnostic training datasets (Ag-MBPP-X, Ag-Codeforces-X, Ag-LiveCodeBench-X), training code, and ready-to-use configurations, making RL training for any programming language as easy as a few lines of YAML.
@@ 1 Introduction
Large language models (LLMs) are remarkably good at programming tasks, especially when coding in high-resource programming languages such as Python and JavaScript. Their proficiency in low-resource programming languages, such as Fortran, Julia, and others, is far more limited. This gap appears both on benchmarks (Cassano et al., 2023) and in popular discourse. Many low-resource languages are adapted to and widely used in particular sectors such as computational science (e.g., Julia, Fortran), medicine (e.g., Mumps), data science (e.g., R), and others. Methods for improving LLMs on such languages would help programmers in these sectors truly take advantage of LLMs.
@@
The capability gap between high-resource and low-resource programming languages occurs for two reasons. First, there is vastly more training data for some languages. For example, The Stack V2 (Lozhkov et al., 2024a), the largest public training corpus of code, has ${\approx}200$GB of Python but only ${\approx}2$GB of Julia and Fortran. Thus pretraining on code makes models significantly better at Python. A subtler reason is the availability of post-training datasets and techniques. Contemporary LLMs are developed with an extensive post-training process that relies on (a) high-quality curated data for supervised fine-tuning, and (b) carefully designed environments for reinforcement learning, which must be able to execute and verify model-generated solutions. Both of these require significant human expertise, which is hard to find for low-resource programming languages.
@@
Our goal in this work is to facilitate post-training LLMs on low-resource programming languages, working towards closing the resource gap. Our key idea is that for a large class of programming tasks, correctness can be stated as a property not of functions or code snippets, but of the entire program’s observable behavior (e.g., I/O). Furthermore, if its correctness can be tested with a verifier program, such a verifier with appropriate problems and test cases can be used to make a universal reinforcement learning environment which can be instantiated for nearly any programming language. In fact, the verifier’s implementation language is independent from the one being learned. This approach matches the formulation of some existing post-training datasets (even if they are intended for Python/C++), and we can reformulate language-specific datasets into this format with LLMs.
@@
Our approach, Agnostics, works based on this insight as follows (Figure 1). 1) We use an LLM to reformulate language-specific datasets into our uniform language-agnostic format. 2) To target a particular language, we generate prompts and instantiate the verifier based on a small (4-5 line) configuration file. 3) We apply reinforcement learning with verified rewards (RLVR) using a robust, language-agnostic execution sandbox that we develop. 4) The result is a model specialized to the target language. Agnostics particularly excels at finetuning models for low-resource languages, as it does not rely on high-quality datasets specific to a particular language.
@@
    Agnostics, a post-training pipeline for coding in arbitrary programming languages;
@@
    The best-performing open-weights ${\leq}16$B models for Lua, R, Julia, OCaml and Fortran;
@@
    Three Agnostics datasets: Ag-MBPP-X, Ag-Codeforces-X, and Ag-LiveCodeBench-X, based on MBPP (Austin et al., 2021), Open-R1 Codeforces (Penedo et al., 2025) and LiveCodeBench (Jain et al., 2024b) respectively.
@@
    A small and carefully designed Agnostics training framework, including a parallel code execution sandbox, sampling, rewards computation, GRPO, and model back-propagation.
Our leaderboard, data, code, and artifacts are referenced at agnostics.abgru.me and in Reproducibility Statement.
Figure 1: Overview: Agnostics Data Preparation and Training. (1) We reformulate existing coding datasets to our format. (2) We adapt the language-agnostic datasets to a particular programming language. (3, 4) We reinforce coding via Group Relative Policy Optimization (Shao et al., 2024; DeepSeek-AI et al., 2025), verifying the programs in our code execution sandbox.
@@ 2 Background and Related Work
Figure 2: High-resource (left) and low-resource (right) languages in the Stack v2.
General-purpose LLMs have been pretrained on code for several years, both because LLMs are widely used for practical programming tasks, and because pretraining on code improves their general reasoning abilities (Ma et al., 2023). Code-specialized models may be trained from a general-purpose model checkpoint or exclusively on code (e.g., Xu et al. (2022); Allal et al. (2023)).
However, the publicly available pretraining data for code is heavily skewed toward a handful of programming languages. E.g., consider The Stack V2 (Lozhkov et al., 2024b), the largest public code pretraining dataset, with code from GitHub and dozens of other sources. The Stack V2 is dominated by relatively few programming languages: just 10 of 619 languages account for over 90% of the dataset. We want to develop models for low-resource programming languages that each account for ${\leq}0.5\%$ of publicly available code (figure 2). We can imagine working around this data scarcity in a few ways. However, previous work shows that up-sampling low-resource languages during pretraining only leads to small benchmark improvements (Orlanski et al., 2023), while fine-tuning on the pre-training data for low-resource languages has negligible impact (Cassano et al., 2024). Thus, it is not clear how to make further gains from existing natural data.
@@
If natural data is not available for a task, it is possible to use LLMs to generate synthetic fine-tuning data (Wang et al., 2023), and there are many techniques for building code-centric supervised fine-tuning datasets (e.g. Luo et al. (2023); Wei et al. (2024c; b)) that work remarkably for Python. Although these approaches could in principle be applied to any programming language, Cassano et al. (2024) show that without distillation or verification (e.g., see Hu et al. (2025); Wei et al. (2024a)), synthetic tasks and code for low-resource programming languages are low-quality, and models fine-tuned on them perform poorly.
MultiPL-T (Cassano et al., 2024), similar to TransCoder-ST (Roziere et al., 2021) and CMTrans (Xie et al., 2024), couples synthetic data generation with verification using rejection sampling: it generates up $n$ candidate programs in a target, low-resource language and only fine-tune models on generations that pass hidden unit tests that it translates from Python. However, the MultiPL-T approach has two significant limitations. (1) For the verifier to not reject all samples, the model must be able to generate a working program within $n$ attempts. In MultiPL-T, ${\approx}30\%$ of prompts produce a working program for $n\in\{50,100\}$ attempts, and the rest are discarded. We train on much harder problems, and estimate that rejection sampling would require an order of magnitude more resources for a comparable acceptance rate (§ 4.3). (2) For each low-resource language of interest, MultiPL-T requires writing a little compiler to translate test cases and function signatures from Python to the target language. MultiPL-T only supports a limited set of built-in Python types (e.g., no classes) and dictates that all Python types and values must faithfully map to the target language. However, depending on the problem and language, the natural data representation may not map cleanly to Python. This can lead to peculiar, unidiomatic translations that require deep language expertise to get right. The Agnostics approach is far easier to use than MultiPL-T, and only requires the user to know how to compile and run a program in the target language from the shell.
@@
DeepSeek R1 (DeepSeek-AI et al., 2025) popularized RL on LLMs with rule-based rewards, instead of learned reward models. R1 reports applying RL to coding tasks without further dataset details. A number of papers apply RL to the NL to code task (Zeng et al., 2025; Gehring et al., 2024; Jain et al., 2025). These techniques target Python and show that RL can improve LLM capabilities beyond what supervised fine-tuning allows alone.
However, the key benefit of RL is that it can train a model to do tasks for which high-quality supervised fine-tuning data is unavailable. There are recent examples of using RL for code optimization (Du et al., 2025; Nichols et al., 2024), resolve GitHub, issue resolution (Wei et al., 2025), and iterative development (Zhou et al., 2025). These papers target tasks in high-resource languages (C++, Java, and Python) whereas Agnostics targets several low-resource languages.
@@ 3 The Agnostics Approach
Our approach comprises (1) a data preparation stage which reformulates language-specific programming tasks to be language-agnostic, and retargets language-agnostic datasets to a programming language of interest (1, 2 in figure 1); and (2) the training stage which uses the GRPO algorithm and an efficient, language-agnostic verification framework (3, 4 in figure 1). Our tasks ask for programs by describing their behavior. The test cases are samples of this behavior, and a verifier program can check if a solution behaves according to the sample. In this paper, we limited ourselves to working with tasks asking for programs which read data from the standard input, compute a unique answer, and write it to the standard output. Hence, the datasets we prepared share one verifier.
# Write a python function to
# identify non-prime numbers.
@@
(a) An MBPP task prompt and associated tests (in gray).
Instruction: Given an integer NN ($N\geq 2$), determine if it is a non-prime number. Output ‘True’ if the number is non-prime, ‘False’ otherwise. Input format: a single integer NN ($N\geq 2$). Output format: a single line containing ‘True’ or ‘False’.
@@ 3.1 Dataset Preparation
Some datasets, like Open-R1 Codeforces (Penedo et al., 2025), already define tasks in the desired I/O style. More commonly, however, code datasets provide a set of unit tests. Figure 3(a) shows a representative item from MBPP: it has a natural language problem description and a Python function signature that comprise the prompt, and a suite of tests used to test model-generated code. These datasets can be easily translated into the I/O format.
To make such problems language-agnostic and compatible with our verifier, we prompt an LLM to reformulate each task so that the program communicates exclusively via plain‑text standard in and standard out. We ask the model to spell out concrete I/O conventions—number of decimal places, newline versus comma separators, ordering of values, and so on—so that the expected behavior is unambiguous. Figure 3(b) shows the reformulated example. § A presents the prompt we use to reformulate MBPP; other datasets might require small changes to the prompt.
@@ 3.2 Programming Language Preparation
To prepare a new language, we author a small configuration file with two purposes. First, it defines a prompt prefix (prepended to each problem by the trainer) which instructs the model to produce code in the target language. Second, the configuration file specifies the shell commands to install the language toolchain and run code. In our experience, a prompt prefix simply asking for a solution in language $L$ is enough for more widespread languages with ${\geq}5\%$ base accuracy. However, when starting from near-zero accuracy, a longer prefix can help prevent common mistakes. E.g., our R language configuration (figure 4) features a longer prompt explaining the quirks of I/O APIs in R. [^1] There are 3 ways to run R, 3 I/O APIs, and only one portable way to read from standard in.
If a model barely knows a programming language, a good prefix can help it. Still, writing the prefix takes manual effort. For OCaml and Fortran, we let a base model generate several faulty snippets, and asked a capable LLM (OpenAI o3) for advice based on the snippets with the following prompt.
We use the resulting instructions verbatim (see § B) when training models. The prefix only needed to slightly raise the model’s train split performance; base accuracy as low as $0.09\%$ was enough for the model to start learning (see § 4.3). Configuring the two languages took 1 hour each.
@@
Use R version 4. Use readLines(con = file("stdin"))‘ to read from
stdin. Use the n argument to read the first n lines. For example:
@@
cat(n) # print the first line of input
Also, please remember to use cat to print output.
Figure 4: An Agnostics configuration snippet for R (slightly rephrased for presentation).
@@ 3.3 Trainer and Code Execution
The Agnostics trainer uses the Group-Relative Policy Optimization (GRPO) reinforcement learning algorithm (Shao et al., 2024), with verifiable rewards (DeepSeek-AI et al., 2025), and further common tweaks to improve its efficiency (Yu et al., 2025). We couple the algorithm with a language-agnostic code execution framework designed to be robust and efficient.
@@
The trainer instantiates the GRPO algorithm as follows. Let $(x,\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K})\sim\mathcal{D}$ be a dataset of language-agnostics tasks, where $x$ is the task prompt and $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$ is the set of I/O examples. Let $P$ be $L.\texttt{prompt}$ from a language configuration $L$ (e.g., figure 4). From the behavior policy $\pi_{\theta_{\text{old}}}$ we sample a group $G$ of candidate responses $\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}({\cdot}\mid{P,x}).$ We assign each candidate a reward $R_{i}$, with $R_{i}=1$ if the execution environment (described later) verifies that the extracted program behaves as in the I/O examples $(\mathit{in}_{k},\mathit{out}_{k})$ and $R_{i}=0$ otherwise. We turn group rewards into sequence-level advantages $\hat{A}_{i}$, and update the policy with the objective
$\mathcal{L}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{\{(x,\_)\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid P,x)\}}$
$\quad\Biggl[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|y_{i}|}\sum_{t=1}^{|y_{i}|}\min\Bigl(\text{clip}\bigl(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon\bigr)\hat{A}_{i},r_{i,t}(\theta)\hat{A}_{i}\Bigr)\Biggr],$
$\text{where}\qquad r_{i,t}(\theta)=\frac{\pi_{\theta}(y_{i,t}\mid P,x,y_{i,<t})}{\pi_{\theta_{\text{old}}}(y_{i,t}\mid P,x,y_{i,<t})},\qquad\hat{A}_{i}=\frac{R_{i}-\operatorname{mean}(\{R_{j}\}_{j=1}^{G})}{\operatorname{std}(\{R_{j}\}_{j=1}^{G})}.$
We omit the KL-divergence term, similar to Yu et al. (2025). We also considered and decided against a reward for a partially-correct answer (see § C.3). We tried to reward the model for code which runs without errors but produces wrong output or for code which only passes the public tests (if there are any). In both cases the models were very likely to learn how to exploit the reward, e.g., by producing empty programs or by hard-coding the public tests (and claiming to produce a “draft answer”).
@@
Our verifier, a language-agnostic code execution sandbox, (1) extracts a program from each candidate; (2) compiles it if needed; and (3) tests it on the I/O examples $\{(\mathit{in}_{k},\mathit{out}_{k})\}_{k=1}^{K}$.
To extract the code, we instruct the model to put it in a Markdown block, which all major instruction-tuned models do by default. Since we rely on the native format of the model, we do not need to train the model with a format reward. This guarantees that the increases in rewards we see during training are real improvements and not merely the result of the model learning to format correctly.
For each language, we build and cache an OCI (2025) container using the configuration $L$. To build the container, we install the language compiler and runtime (the script $L.\texttt{install}$), and include a generic execution harness which runs and tests candidate programs. The execution harness runs continuously in the container, waiting for triples with the candidate program, the set of input/output examples, and timeouts. The harness (1) writes the program to disk (to $L.\texttt{filename}$), (2) compiles it if needed ($L.\texttt{compile}$), (3) runs it on each received input ($L.\texttt{execute}$), and verified that it produces the expected output. The harness imposes timeouts on the compilation step and each execution, and returns reward 0 on any timeout or failed verification. It is important to have timeouts for both compilation and execution. This prevents pathologies such as unbounded macro expansion in Julia (caught by the compile timeout) and infinite loops (caught by the execution timeout). Using containers also allows us to limit CPU, memory, and filesystem usage; no elevated privileges are granted to the generated program. Although the current datasets only specify tasks by standard I/O, the same sandbox can safely accommodate problems that read/write from network or disk.
@@
Finally, to improve compile times, our execution environment mounts a RAM disk in each container. Compilation may be slow due to creating many intermediate files, and indeed some large C++ projects, e.g., Firefox, recommend using a RAM disk to speed up their builds (Firefox, 2025).
@@
We implement the trainer and execution environment with Ray (Moritz et al., 2018), which facilitates multiprocessing and distributed computing. In particular, Ray lets us distribute the training over a network of heterogeneous nodes, which allows running the trainer on a node specialized for GPU work and the execution environment on a node specialized for CPU work. Ray also lets us easily separate group generation and loss computation into inter-communicating processes. Running the two in parallel significantly speeds up training, as we found that they take a roughly comparable amount of time. The execution environment is also an actor and manages containers with Python asyncio coroutines, not actors, to minimize inter-process data copying.
@@
We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a learning rate of $5\times 10^{-6}$ and a cosine decay schedule with a warmup of 0.1 epochs. We process 4 prompts in each batch, with group size 32 per prompt. When training we use temperature 0.7 and disable reasoning in hybrid models. (Many generations still show reasoning-like text, either before the answer or in comments.) We describe how we chose the hyperparameter values and how we evaluted different values in § C.1.